OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1

OpenStudy (anonymous):

(3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

Plug in the equation and it shows you the steps too.. http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

I still don't get it

OpenStudy (anonymous):

Like, it gives me an equation.

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (anonymous):

−2x+y+2x=1+2x(Add 2x to both sides) y=2x+1

OpenStudy (anonymous):

(2x+1) for y in −4x+y=−1:

OpenStudy (anonymous):

enter them in here and it should give you all the info you need, i use this all the time.

OpenStudy (anonymous):

−4x+2x+1=−1 Add -1 to both sides)

OpenStudy (anonymous):

−2x=−2 (Divide both sides by -2)

OpenStudy (anonymous):

Substitute (1) for x in y=2x+1:

OpenStudy (anonymous):

y=2(1)+1